Use Gauss-Jordan Elimination to solve the following system of equations.
2x + 2y 6z = −2
x + 5y + z = −3
6x + 14y 10z = −8
Answer #1

solution Giren: System of Equations → - 2x + ay 62= -2 x+ sy +2=-3 6x + 4y - 10 z=18 (to rolre using the Gaurs Jorden eliminaform, so, converting Rio R/ to now reduced echelon r1 1 - 3 1 -17 11 5 11 -3 6 14 -10 1-8) - 1 To T R₂ R₂ - R, ✓ R3 R₃-6R, 1